# Optimal swimming with body compliance in an overdamped medium

### v1

@@ Abstract
Elongate animals and robots use undulatory body waves to locomote through diverse environments. Geometric mechanics provides a framework to model and optimize such systems in highly damped environments, connecting a prescribed shape change pattern (gait) with locomotion displacement. However, existing approaches assume precise execution of prescribed gaits, whereas in practice environmental interactions with compliant bodies of animals or robots frequently perturb the realized trajectories. In this work, we extend geometric mechanics to predict locomotor performance and search for optimal swimming strategy of compliant undulators. We introduce a compliant extension of Purcell’s three-link swimmer by incorporating series-connected springs at the joints. Body dynamics are derived with resistive force theory. Geometric mechanics is incorporated into movement prediction and into an optimization framework that identifies strategies for controlling compliant swimmers to achieve maximal displacement. We validate our framework on a physical cable-driven three-link limbless robot, and demonstrate accurate prediction and optimization of locomotor performance under varied programmed, state-dependent compliance in a granular medium. Our results establish a systematic physics-based approach for modeling and controlling compliant swimming locomotion, highlighting compliance as a design feature that can be exploited for robust movement in homogeneous and heterogeneous environments.
@@ I Introduction
Undulatory locomotion, in which traveling body waves generate self-propulsion, is a fundamental strategy employed by elongate animals (e.g., snakes [1, 2, 3, 4, 5] eels [6, 7], skinks [8, 9, 10]) and robots [11, 12, 13, 14, 15, 16] to traverse both fluid and terrestrial environments. These locomotive capabilities emerge from the interplay between body dynamics and environmental media [17, 18], motivating extensive research efforts aimed at uncovering substantial principles and deriving unified frameworks for understanding animal movements and informing robotic developments [17, 19, 8, 20, 16, 21].
@@
Recent studies focusing on highly damped environments (e.g., viscous fluids [22, 23, 24], granular media [25, 26], and frictional substrates [27, 28]), where inertial effects can be neglected, have shown that undulatory locomotion can be effectively modeled within a general geometric framework known as geometric mechanics (also referred to as the geometric phase approach) [19, 20]. The concept of geometric phase was originally developed in quantum mechanics [29] and later applied to classical mechanics [30]. Specifically, Shapere and Wilczek demonstrated that the geometric phase approach could provide important insights into the locomotion of microorganisms [31]. Further, beginning in the 1980s, physicists and control theorists extended the geometric phase framework to a variety of systems moving through different environments. For example, it has been applied to Purcell’s three-link swimmer in viscous fluids [32, 33] and in dry granular media [19], where the dynamics can be determined analytically and empirically. More broadly, with appropriate dimension-reduction techniques, geometric mechanics has proven effective in modeling animal locomotion across diverse environments [20] and in guiding the control of limbless robots [16, 34].
@@
Though effective in modeling undulatory locomotion in highly damped environments, geometric mechanics tools are limited to linking determined self-deformation patterns (gaits) with their resulting locomotor performance [31]. However, in the practical control of limbless robots or the neuromechanics of elongate organisms, shape-change patterns are not directly prescribed but instead emerge from interactions between body viscoelasticity and the surrounding homogeneous or heterogeneous environment. This interaction-involved “loop” plays a central role in the gait adaptation of C. elegans [35, 36, 37] (Fig. 1(c)) and has been shown to simplify robot control in rheologically complex media [38]. To address this, Ramasamy and Hatton [39] introduced preliminary insights by modeling Purcell’s three-link swimmer with one actuated joint and one purely elastic joint. Their geometric framework successfully predicted performance in such compliant, underactuated systems, but it did not directly support gait optimization, and its broader applicability to physical robotic platforms remains to be demonstrated.
@@
Similar attention has been devoted in robotics to the optimal control of systems with elastic actuation [40, 41, 42, 43, 44]. Inspired by the compliance observed in biological organisms, elastic actuation—achieved through actuators incorporating elastic components or soft materials—has been widely implemented in manipulators [45, 46, 47], legged robots [48, 49, 50, 51], and exoskeletons [52, 53, 54] to mitigate external perturbations, for safer physical interactions and/or improved adaptability to complex environments. However, the added compliance complicates control by introducing nonlinear dynamics and frequency-dependent behavior, particularly in contact-rich tasks. As a result, controlling such systems often requires computationally intensive methods, which can limit their practical deployment in real-world applications [55, 45, 56, 57].
@@
In this work, we extend geometric mechanics to model compliant swimmers, in which the realized gait emerges from interactions between body compliance and the environment. Further, we develop a general optimization framework for maximizing displacement in compliant swimmers across a wide range of highly damped environments, including viscous fluids, granular media, and frictional substrates. To achieve this, we extend Purcell’s three-link swimmer to a compliant form by modeling each motor as connected in series with a torsional spring. We first formulate the system dynamics and derive the associated height function to predict locomotor performance. We then propose a gait optimization framework that identifies gaits that yield maximal displacement from height functions and generates their corresponding motor commands through inverse body dynamics. We validate the framework by applying it to a physical cable-driven limbless robot capable of exhibiting tunable, state-dependent compliance (Fig. 1(a) and (b)) and demonstrate that it can accurately predict and optimize its locomotion under varied body compliance in a granular medium.
@@ Acknowledgements.
The authors are grateful for funding from Army Research Office Grants W911NF-11-1-0514 and W911NF-20-1-0129. The authors acknowledge the use of ChatGPT (OpenAI) for language editing.

### v2

@@ Abstract
Elongate animals and robots use undulatory body waves to locomote through diverse environments. Geometric mechanics provides a framework to model and optimize such systems in highly damped environments, connecting a prescribed shape change pattern (gait) with locomotion displacement. However, the practical applicability of controlling compliant physical robots remains to be demonstrated. In this work, we develop a framework based on geometric mechanics to predict locomotor performance and search for optimal swimming strategies of compliant swimmers. We introduce a compliant extension of Purcell’s three-link swimmer by incorporating series-connected springs at the joints. Body dynamics are derived using resistive force theory. Geometric mechanics is incorporated into movement prediction and into an optimization framework that identifies strategies for controlling compliant swimmers to achieve maximal displacement. We validate our framework on a physical cable-driven three-link limbless robot and demonstrate accurate prediction and optimization of locomotor performance under varied programmed, state-dependent compliance in a granular medium. Our results establish a systematic, physics-based approach for modeling and controlling compliant swimming locomotion, highlighting compliance as a design feature that can be exploited for robust movement in both homogeneous and heterogeneous environments.
@@ I Introduction
Undulatory locomotion, in which traveling body waves generate self-propulsion, is a fundamental strategy employed by elongate animals (e.g., snakes [1, 2, 3, 4, 5] eels [6, 7], skinks [8, 9]) and robots [10, 11, 12, 13, 14, 15, 16] to traverse both fluid and terrestrial environments. These locomotive capabilities emerge from the interplay between body dynamics and environmental media [17, 18], motivating extensive research efforts aimed at uncovering substantial principles and deriving unified frameworks for understanding animal movements and informing robotic developments [17, 19, 8, 20, 15, 21].
@@
Recent studies focusing on highly damped environments (e.g., viscous fluids [22], granular media [23, 8], and frictional substrates [24]), where inertial effects can be neglected, have shown that undulatory locomotion can be effectively modeled within a general geometric framework known as geometric mechanics (also referred to as the geometric phase approach) [19, 20]. The concept of geometric phase was originally developed in quantum mechanics [25] and later applied to classical mechanics [26]. Specifically, Shapere and Wilczek demonstrated that the geometric phase approach could provide important insights into the locomotion of microorganisms [27]. Further, beginning in the 1980s, physicists and control theorists extended the geometric phase framework to a variety of systems moving through different environments. For example, beginning with Purcell’s three-link swimmer [28], a simple model used to study locomotion in low–Reynolds-number fluids [29, 30], geometric mechanics has been extended to viscous fluids [31, 32] and dry granular media [19], where the dynamics can be determined analytically and empirically. More broadly, with appropriate dimension-reduction techniques, geometric mechanics has proven effective in modeling animal locomotion across diverse environments [20] and in guiding the control of limbless robots [15, 33, 32].
@@
Though effective in modeling undulatory locomotion in highly damped environments, geometric mechanics tools are limited to linking determined self-deformation patterns (gaits) with their resulting locomotor performance [27]. However, in the practical control of limbless robots or the neuromechanics of elongate organisms, shape-change patterns are not directly prescribed but instead emerge from interactions between body viscoelasticity and the surrounding homogeneous or heterogeneous environment. This interaction-involved “loop” plays a central role in the gait adaptation of C. elegans [34, 35] (Fig. 1(c)) and has been shown to simplify robot control in rheologically complex media [36]. To address this, Ramasamy and Hatton [37] extended the geometric framework to identify optimal gaits for compliant swimmers. The framework successfully modeled and optimized performance under system constraints for two types of proposed Purcell’s three-link swimmers, each with one actuated joint and one purely elastic joint, in a viscous fluid (similar to the systems described in [38]). However, its broader applicability to physical robotic platforms remains to be demonstrated.
@@
Similar attention has been devoted in robotics to the optimal control of systems with elastic actuation [39, 40, 41, 42, 43]. Inspired by the compliance observed in biological organisms, elastic actuation—achieved through actuators incorporating elastic components or soft materials—has been widely implemented in manipulators [44, 45, 46], legged robots [47, 48, 49, 50], and exoskeletons [51, 52, 53] to mitigate external perturbations, for safer physical interactions and/or improved adaptability to complex environments. However, the added compliance complicates control by introducing nonlinear dynamics and frequency-dependent behavior, particularly in contact-rich tasks. As a result, controlling such systems often requires computationally intensive methods, which can limit their practical deployment in real-world applications [54, 55, 56].
@@
In this work, we develop a framework based on geometric mechanics to model and optimize the performance of compliant swimmers in overdamped environments. To achieve this, we introduce a compliant extension of Purcell’s three-link swimmer by incorporating series-connected springs at both joints (as shown in Fig. 2), in which the realized gait emerges from interactions between actuation patterns, body compliance and the environment. We first model the system dynamics using resistive force theory (RFT) and employ geometric mechanics tools to predict locomotor performance. We then propose a gait optimization framework that identifies gaits that yield maximal displacement from geometric mechanics and generates their corresponding control commands through inverse body dynamics. We validate the framework by applying it to a physical cable-driven limbless robot capable of exhibiting tunable, state-dependent compliance (Fig. 1(a) and (b)) and demonstrate that it can accurately predict and optimize its locomotion under varied body compliance in a granular medium.
@@ Acknowledgements.
The authors thank Ross L. Hatton for helpful discussion. The authors are grateful for funding from Army Research Office Grants W911NF-11-1-0514 and W911NF-20-1-0129. The authors acknowledge the use of ChatGPT (OpenAI) for language editing.